### v7

# Analytic 3D vector non-uniform Fourier crystal optics in arbitrary dielectric

Abstract
To find a suitable framework for nonlinear crystal optics(NCO), we have revisited linear crystal optics(LCO). At the methodological level, three widely used plane wave bases are compared in terms of eigenanalysis in reciprocal space and light field propagation in real space. Inspired by complex ray tracing, we expand M.V. Berry and M.R. Dennis’s 2003 uniform plane wave model to non-uniform Fourier crystal optics and ultimately derive the explicit form of its 32 transition matrix, bridging the two major branches of crystal optics in reciprocal space, where either ray direction or spatial frequency serves as the input variable. Using this model, we create the material-matrix tetrahedral compass to conduct a detailed analysis of how the four fundamental characteristics of materials (linear/circular birefringence/dichroism) influence the eigensystems of the vector electric field in two-dimensional spatial frequency domain and its distribution in three-dimensional space with a crystal-2f configuration. Along this journey, we have uncovered new territories in LCO in both real and reciprocal space, such as infinite singularities arranged in disk-, ring-, and crescent-like shapes, “L shorelines” resembling hearts, generalized haunting theorem, double conical refraction, and optical knots it induces. We also present our model’s early applications in focal engineering and NCO. As the opening chapter in a trilogy, this work interweaves crystal optics, Fourier optics, and nonlinear optics, while integrating theoretical, computational, and experimental physics, advancing all six domains.
keywords:
Crystal optics, Singular optics, Fourier optics, Nonlinear optics, Quantum opticsContents
1 Introduction
Threefold purpose guides us: to establish crystal optics(CO) and Fourier optics(FO) as the physical and computational bases of nonlinear crystal optics(NCO) in the first part of a trilogy; to propose that three plane-wave bases form the mathematical core of CO spectral analysis, each offering trade-offs; to seek and knock the computational boundaries of linear crystal optics(LCO) and NCO via large-scale FO simulations in both three-dimensional(3D) reciprocal and two-dimensional(2D) spatial domains.
Out of the 32 macroscopic crystal symmetries, 21 display non-centrosymmetry, enabling second harmonic generation(SHG) within these point groups. However, among these 21 crystal classes, 18 are birefringent or anisotropic, and more than half() exhibit natural optical rotation, demonstrating intrinsic optical activity(OA) or chirality[1]. As a result, NCO often encountered tremendous computational challenges[2, 3, 4], due to the complexity of linear crystal optical(LCO) processes in low-symmetry materials with complex dielectric tensor [5].
Think about the ordinary(o) and exordinary(e) waves in nonlinear crystal optical(NCO) phase-matching types[6, 7, 8, 9, 10, 11, 12](e.g. type-I ooe in Fig. 5a): what are they? Eigenmodes governed by CO[13, 4, 14, 15, 16, 17, 18, 19]!! That is to say, the entire NCO process is constantly subject to the constrains imposed by CO[13, 5, 3, 4], which is itself inherently complex[20]. More phenomenologically, for all frequency-mixing phenomena, both the pumps/fundamental waves(FWs) and the newly generated frequencies(NFs) within the crystal, must be first decomposed into the material’s intrinsic eigenmodes, then diffract independently (and anisotropically) during subsequently linearly superposition and interference, while potentially enduring absorption or gain[21, 22].
Beyond the anisotropy of the uniformly distributed itself, the micro-nano structures modulated by fabrications on the material’s and [23, 24, 25, 26, 27, 28, 29, 30] tensors in recent engineering developments, likewise call for an immediate and precise consideration of the linear optical(LO) scattering[31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 27] and diffraction[41, 42, 43, 44, 45, 29, 30] effects caused by the material’s (sub)wavelength-scale inhomogeneity, both for FWs[46, 47, 48, 33, 39] and NFs [49, 22, 50, 51, 32, 40, 27, 28, 29, 30].
Hence, from first principles, NCO is fundamentally built upon LCO[13, 3, 4], which in turn is founded on linear optics(LO) and CO. This implies that NCO simultaneously inherits all the challenges arising from both CO and LO, each of which has undergone an extraordinarily convoluted path of mathematical development spanning nearly two centuries — for LO, from ray optics[52, 53, 18, 54, 55] to diffraction integrals[56, 57, 58, 59, 60, 61] and eventually to FO[62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]; for CO, from uniform plane-wave models[13, 4, 74, 75, 76, 77, 78, 79, 80, 81, 1, 5, 20] to non-uniform plane waves [82, 83, 84, 15, 85, 14, 86, 87], [88, 89, 90, 91, 92, 93] and finally to matrix exponentials [94, 95, 96, 97, 98, 92].
Given that the impact of LCO(= LO CO) on NCO has been historically either ignored or modeled incorrectly[22], we undertake a thorough reexamination of LCO in Supplementary Note 1. Through this inquiry, we summarize that:
The computational boundary of LCO, and thus of NCO, terminates at optical singularities, i.e. exceptional points(EPs) in 2D reciprocal space. The optical field along such singular directions is, in principle, uncomputable due to the degeneracy/parallelism of paired eigen-polarization states when using spectral methods like plane wave based FO.
While matrix exponential[95, 96, 97, 98] do provide a workaround and Jordan decomposition[99, 100, 99, 101, 102, 97, 103, 104, 105, 106] further allows for an internal probe of these singularities[94, 107, 108, 109], the latter becomes invalid when applying large-scale Fourier optical(FO) sampling[97, 99], and both suffer from other inherent computational limitations[96, 92, 97, 110] (see Supplementary Note 1). As a result, we avoid employing matrix exponentials as the crystal optical(CO) Fourier basis in this LCO model for massive numerical experiments.
Only two options remain (see “Methods”): uniform CO basis that contradicts both FO and boundary conditions(BCs), but only requires solving bi-quadratic equations thanks to spherical coordinates()[5]; rectangular() non-uniform CO basis that fits both physics(FO) and experiments(BCs), but demands batchly solving quartics with intractably long formulas[96].
In this Article, we propose a non-uniform linear Fourier crystal optical(LFCO) model, whose CO part integrates the advantages of both uniform[75, 76, 77, 78, 79, 80, 81, 1, 5, 20] and non-uniform[83, 15, 85, 14, 84, 86, 87, 111, 112] plane wave models , solving a pure bi-quadratic under FO and boundary constraints. This LFCO model extends the standard non-diagonal/non-positive-definite[113, 114, 115, 116, 117, 118, 119, 113, 120, 121, 122]/non-Hermitian[123, 124, 125, 126, 127, 128, 129, 130, 131, 132]/non-unitary[133, 134, 135, 136]/non-normal[107]/non-(mirror-[137])symmetric[5, 138, 139]/non-reciprocal[140, 141, 137] 22[142] transfer[98, 143]/transition[98, 110] matrix( evolution operator[144]) in CO to a 32 form (see Eqs. 5 and 6), enabling explicit fully vectorial[93, 91, 145, 146, 147, 148, 149, 150, 151] computation within arbitrary materials, and at the same time, allows the LO part to directly operate in the non-uniform[83, 15, 85, 14, 84, 86, 87, 92, 84] FO framework[91, 62, 90, 93, 89].
To indirectly validate this LFCO model, and highlight its powerful role in advancing the analytic development of NCO, we offer a roadmap for expanding the LFCO model in Fig. 2c. Following its conceptual architectures — scalar(, semi-vector,) and (full-)vector NFCO — we first demonstrate this LFCO model’s applications: in full-vector NFCO, via chiral second-harmonic conical refraction(SHCR) in Figs. 1 and 2, and harmonic spin-orbit(S-O) angular momentum(AM) cascade in Fig. 4 (see also Figs. SC3-SC5); in scalar NFCO, through phase-matching-controlled orbital angular momentum(OAM) conversion in Fig. 3 (see also Figs. SB2-SB6); and in semi-vector NFCO, via full conical phase matching(FCPM) in Fig. 5 (see also Fig. SC2).
To directly validate this LFCO model, as stated at the beginning of Section 2.4, the procedure must begin with verifying the correctness of the eigenvalue-eigenvector computations (Fig. S8), followed by confirming the accuracy of the field distributions in real space (Fig. 7, Fig. 8, Figs. S9-S11, and Figs. SA2-SA10).
Accordingly, we introduce the material-matrix(M-M) tetrahedron compass(TC) in Fig. S8, where scanning parameters along three of its edges depict a theoretical panorama of the adiabatic evolution(AE) of electric field eigenmodes in the domain[152], competing among the three primary material properties, i.e. linear dichroism(LD), circular dichroism(CD), and optical activity(OA). During this process, we observe that CD, adhering to the haunting (C points) theorem[5] as well, can lead to heart-shaped L shorelines (as an upgrade to the L lines in CO[5]) and an infinite array of singularities in 2D domain, often arranged in patterns resembling disks, rings[101, 153], or crescents (instead of 8 finite EPs[5, 154, 154, 155, 22, 92, 156]). In Fig. S8, certain known phenomena reappeared[101, 153, 157], while the others, to the best of our knowledge, are originally predicted.
Aided by crystal-2f system proposed from Fig. S9, we map out a second, more experimentally relevant[158, 159, 139, 81, 160] panorama in the 3D domain in Fig. 7, freezing the AE of the light field distribution between the three vertices of the M-M TC, corresponding to the material’s birefringence(Bi), LD, and OA.
Still in real space, but with higher numerical aperture(N.A.), we offer a unified solution for the forward propagation and inverse design of focal fields in Fig. 8, where a dual-eigenmode decomposition is proposed to explain Raman spikes[161] in conical refraction(CR) arising solely from slow modes, multifoci induced by laser processing inside materials, and aberration correction without Zernike polynomials. Along the way, we have observed two surprising new phenomena: double conical refraction(DCR) and the resulting optical field knots.
Moreover, our trilogy forecasts the future of quantum crystal optics(QCO) quantum optics(QO) by demonstrating spontaneous parametric down-conversion(SPDC) in 3D nonlinear photonic crystals(NPCs), beyond the extensive faithful reproduction of past sophisticated experimental results in both LCO and NCO.
This work unites mathematical elegance(see Supplementary Notes 2-5), physical intuition(see Supplementary Notes 7), and experimental accuracy(see Supplementary Note 8-9). Even so, it too currently suffers computational limits(see Supplementary Notes 5-6).
2 Results
2.1 By 2025, it still remains a great challenge to reproduce chiral SHCR
To establish the technical soundness of this LFCO approach, we first present results from the third-stage development — beyond the scope of, yet based on the LFCO model itself — namely, full-vector nonlinear Fourier crystal optics(NFCO) simulations versus experiments.
The optically active( chiral[160]) SHCR[159, 162, 163, 164, 165] experiment in Fig. 1f[164] represents an NCO experiment that, in principle, cannot be well-reproduced without a mature LCO framework[165, 166, 167, 168, 163, 169, 170, 171, 164, 172, 173]. Almost all previous NCO modeling attempts[170, 168, 168, 163, 167] have invariably failed to capture the experimental phenomena at a pixel-level resolution far beyond phenomenology. Other works refrain from modeling altogether[164, 165, 171].
As the concluding chapter[163] of Bloembergen’s trilogy[159, 162, 163], SHCR is nothing less than a Holy-Grail level[174] modeling challenge as an open benchmark in NCO computation, demanding equally deep command of LCO and NCO alike. The absence of either renders the process inherently intractable.
2.2 The union of CO, NO, and FO: a theoretical, experimental, and computational Holy Grail
The built-in difficulty of (chiral-dichroic) SHCR simulation is discussed through four lenses — LCO, NCO, phenomenology, and computation. As to its mathematical/theoretical/modeling challenges, for now, must be deferred, to the latter two parts, of the trilogy.
On a phenomenological level, this NCO process(i.e. SHCR) involves synchronous but distinct LCO anisotropic diffraction for both the fundamental and harmonic waves(HWs). In Fig. 1f, the FW() conical refracts(CR) along its optic axis, while the HW() tends to double refract(DR) (yet mixed with CR, see Fig. S9 in Supplementary Note 8), forming a CRωCR-DR[134])2ω combination. In Fig. 2, pumped along the harmonic’s axis, the configuration shifts to DRωDR-CR[175])2ω. Each Fourier component of each eigenmode, in both 2D spatial and one-dimensional(1D) angular frequency domains , exhibits wave vector double(DR)/conical(CR)/double-conical(DR-CR) or even double conical refraction(DCR) (see Fig. 8d) at the material interfaces, superimposed on walk-off between Poynting vector and wave vector inside material.
Challenge deepens: what if the material is also chiral[144, 165, 164, 176, 160, 133, 177] and dichroic[134, 21, 22, 160, 139, 178, 157, 81], beyond being birefringent[22, 179, 134]? The corresponding LCO-based NCO process lies at the pinnacle of linear crystal optics(LCO), a domain shaped over nearly 200 years by the world’s leading LCO theorists[94, 180, 92, 181, 182, 183, 181, 184, 75, 185, 186, 187, 188, 189, 190, 191, 192, 5, 20, 101, 193, 194, 195], experimentalists[158, 162, 159, 196, 133, 197, 198, 199, 200, 201, 181, 202, 139, 178, 157, 81, 203, 204], applied[205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229], and computational[230, 95, 231, 46, 94, 96, 98, 232, 90, 211, 44, 233, 234, 235, 236, 237, 238, 54, 239, 240, 241, 62, 89, 91, 242, 243, 244] physicists.
Challenge deepens twice again: above LCO processes repeat itself, acting concurrently and independently across all wavelength components — discrete for continuous-wave pumping, continuous for ultrafast excitation. Consider SHG, the hallmark of NCO: fundamental diffraction( LCO), harmonic diffraction( LCO), up-conversion( NCO), and down-conversion(energy backflow NCO) — each persists in isolation, requiring none of the other three. These 4 dynamics are not stepwise[41], not sequential[245, 246], not cascading[247, 194, 248, 249, 207, 250, 251, 252, 253], but fully simultaneous[254, 255, 250, 247, 256, 257, 258, 237]. From the very first moment the pump touches the crystal, all four unfold everywhere, at every instant, until the harmonics exit through both end faces, or the pumping stops.
So, we ask, how, exactly, to calculate the Chiral SHCR process in Fig. 1? Most previous efforts have ignored the foundation of NCO — LCO, and unceasingly phenomenologically to decouple LCO(anisotropic diffraction) from NCO(frequency conversion processes), both of which occur in parallel across all spatiotemporal spectral components, all eigenmodes, and all tensor & vector components. Some ask, parallel computation? Yes, but still either mathematically incorrect or numerically inefficient.
On a computational level, each of the six NCO phase-matching types() in Fig. 1a,d (and Fig. 2b) is, for chiral SHCR, phase-mismatched and “very likely” non-degenerate. Wherein, phase mismatch implies that none of these processes dominates the NCO in terms of energy conversion efficiency, necessitating the simultaneous consideration of all six. Non-degeneracy means that outcomes of the six (eigen)mode combinations() are usually distinct(see Fig. 1d), requiring separate calculations for each.
Except for computing all phase-matching types (6 for KTP), a (full-)vector, (phase-)mismatched NCO process also involves all non-zero second-order (nonlinear coefficient) tensor elements (5 for KTP), and all nonlinear source terms (7 for KTP) from pairwise products of the fundamental wave(FW)s’ eigen-polarizations’ x, y, z components. — The ( o, e for KTP) eigen-polarizations and x, y, z Cartesian components are independent, leading to multiplicative complexity.
For KTP, one vector NFCO process takes scalar NFCO runs (Fig. 1c). If one further scans wavelength (pulse injection, see Figs. 4, 1, 2 and 3 and SC2), propagation distance (dynamical evolution, see Figs. S9-S11, Fig. 8 and Fig. 1c), pump power, beam waist (see Fig. S11c), incident angle (see Fig. 2, Fig. 3, and Fig. S9), N.A. (see Fig. 8 and Fig. S11c), temperature , external magnetic field , -frame orientation (see Figs. SA4-SA5), or material coefficients (for adiabatic tuning, see Fig. 7 and Fig. S8) — each adds a for-loop layer, and each layer may grow due to low crystallographic symmetry, or dense parameter sampling.
Therefore, to comprehensively investigate the high-dimensional parameter space of light-matter interactions, as done by Berry et al.[5, 20], Mcleod et al.[91], A.Favaro et al.[187], Hehl et al.[260], this study, and many others[184, 261, 262, 263, 195, 174, 152], both linear and nonlinear crystal optics(L/NCO) call for theories/models/algorithms with a high speed-accuracy product.
The incoming second part of this trilogy (our scalar NFCO model) points out: with a : scale mismatch — crystal macro(10 mm[164]), lightwave meso(1 m) — Green-function formalism[264, 265] struggles transversely/in-plane(x-y), while split-step Fourier[266, 267, 268, 269, 270, 271] and pseudo-spectral method[272, 273, 274] (with Runge-Kutta scheme[275, 44, 276, 253]) strain longitudinally/out-of-plane(z), making forward computation of even one scalar NFCO hard, let alone all 42 in a full-vector NFCO. To thoroughly resolve the scalar NFCO process, we formulate the imminent Nonlinear Angular Spectrum Theory(NLAST), with its validity demonstrated in Fig. 3 (and Figs. SB2-SB6).
All forms of N(F)CO modeling — scalar (Fig. 3 and Figs. SB2-SB6), semi-vector (Fig. 5 and Figs. SC2), or full-vector (Figs. 1, 2 and 4 and Figs. SC3-SC5) — depend on a vector L(F)CO model (Figs. 6, 7 and 8, Figs. S8-S11, and Figs. SA2-SA10) to yield the requisite eigenvectors (eigen-polarizations) and eigenvalues (eigen-wavevector-z-components for FO) , necessary for evaluating the initial complex amplitudes (i.e. modal coefficients) in Eq. (S56c) to calculate frequency-mixing dynamics, and for determining the nonlinear conversion efficiency “eigenvalue mask” sinc “eigenvector mask” in Fig. 5.
To showcase the basics of semi-vector NFCO, Fig. 5j reproduces the full conical phase matching(FCPM) second harmonic generation(SHG) proposed by Belyi et al.[221], illustrating the calculation of conversion in a uniaxial BBO crystal through Fig. 5a-i, where both the first- and second-order susceptibilities are anisotropic.
In subsequent (both scalar and vector) NFCO, transverse wave vector conservation (Fig. 5a,b) is elevated to the same fundamental level as conservation and phase continuity at the boundaries, making it a prerequisite that must take priority. The conversion efficiency, determined solely by the longitudinal phase mismatch (Fig. 5f) and effective nonlinear coefficient (Fig. 5i)[277, 13, 3, 4]
is subordinated to the fulfillment of transverse momentum conservation. Notably, , , and in Eq. 1 represents the eigen-polarization states of , , and in the frame, whose combination are fixed under certain phase-matching type. For example, phase-matching type-I ooe from Fig. 5g,h needs to compute = from Eq. 1. Above -frame eigenvectors are almost directly provided by Berry & Dennis’s 2003 model[5] in Fig. 10b, though a few minor adjustments in Fig. 10c are still required.
As two masks determining the angular distribution of frequency conversion efficiency, the degree of longitudinal phase matching sinc from Fig. 5c and the effective nonlinear coefficient from Fig. 5i, respectively, depend on the eigenvalues and eigenvectors in linear Fourier crystal optics(LFCO), reaffirming that NCO is fundamentally grounded in LCO, and both linear(LO) and nonlinear optics(NO) should ultimately be incorporated into the framework of Fourier optics(FO).
2.3 NO’s two pillars: CO (32 transition matrix) and FO (FT pairs OTFs)
Our LFCO story began with a fleeting glimpse of CO: plugging a plane-wave trial for the electric vector field into the monochromatic wave equation of purely electro-anisotropic medium yields a characteristic equation[91] (see Supplementary Note 2)
for both coordinate-free[74, 278] wave eigenvector field(as FO plane-wave ensembles) and electric eigenvector field , with no explicit independent/free/input variable (ray direction) or (spatial frequency in FO), admitting both uniform spherical solution and non-uniform Cartesian solution . Should be expressed in spherical() coordinates, i.e. , or Cartesian() coordinates, i.e. ?
As discussed, these two Fourier eigenbases each have their strengths and limitations (see “Introduction”, “Method”, and Supplementary Note 1,2): -based CO is a “black box” — for lacking explicitness; -based CO is white-boxed, yet fails FO and BCs.
In order to directly utilize the closed-form eigenvalues, i.e. refractive index , to the bi-quadratics from existing uniform-plane-wave CO models[74, 75, 76, 77, 78, 79, 80, 81, 1, 5, 20], we express the real wave vector direction as a function of spatial frequency as follows
which is a transcendental equation of with being provided by various even-spectrum( biquadratic) LCO models[74, 75, 76, 77, 78, 79, 80, 81, 1, 5, 20]. Eq. 3 can be solved through direct iteration (see Method), Newton’s iteration[241], or other methods (see Discussion and Supplementary Note 6.2), with convergence typically occurring within two iterations when anisotropy is weak and the N.A. is small.
The eigenvalues and eigenvectors (i.e., eigen-polarization states) of non-uniform LFCO can be obtained by respectively substituting obtained through Eq. 3 into
and in Method. Then, as depicted in Fig. 6 and Method, the eigenvalue pairs construct the propagation matrix , while the eigenvector pairs form the eigen-polarization state matrices and . A sandwich multiplication of these three matrices yields the 32 transition matrix (see Supplementary Note 5)
which structurally resembles the matrix exponential after Jordan decomposition (see Supplementary Note 1), realizing the transition from the two-component vector field on the input plane at to the three-component vector field on the output plane at through . Coupled with Fourier transform(FT) pairs (defined in Eqs. (S4, S5)) , Eq. 5 allows us to study the evolution and 3D distribution of the vector electric field in arbitary anisotropic dielectrics in real space, as demonstrated in Fig. 6.
Apart from CO’s in-crystal 32 transition matrix in Eq. 5 and FO’s FT pairs in Fig. 6, this LFCO model and the full trilogy, make broad use of self-built out-of-crystal optical transfer functions(OTFs) for lenses (Figs. 1, 2, 4, 3, 7 and 8 and Fig. S9), objectives (Figs. 8 and 4), half- and quarter-wave plates (Fig. 4), q-plates (Fig. 4), linear polarizers/analyzers (Fig. S11), etc., all within a custom FO framework.
2.4 Material- tetrahedron compass Crystal-2f setup: guidances for batch LFCO numerical experiments in 2D reciprocal and 3D real space
For this LFCO spectral method, the investigative and computational agenda unfolds according to the following sequence: different optical materials different tensors different eigenmodes/eigensystems/eigen-value-vector-pairs different 32 transition matrices different in-crystal field distributions different out-of-crystal field evolution .
The field itself varies over space (2D reciprocal space, for eigensystems in Fig. S8; 3D real space, for wave diffraction in Fig. 7). This work first explores the field’s adiabatic evolution(AE) (Fig. S8 and Fig. 7) or dynamic evolution (Fig. S9 and Fig. 8) under parameter shifts, using the proposed M-M TC( M- TC in this work), implying a “change of the change”.
Arbitrary complex dielectric tensors are typically associated with birefringent-chiral-dichroic dielectrics[153, 279, 280]. In order to relate the mathematical properties of the matrix to the optical/physical properties of the material, we propose the M-M TC in Fig. 7 and Fig. S8.
The M-M TC consists of 6 edges: Hermitian(H), anti-Hermitian(!H), symmetric(S), anti-symmetric(!S), real part(Re), and imaginary part(Im) — corresponding to the six mathematical attributes of a complex matrix ().
The M-M TC consists of 4 vertices: birefringence(Bi), optical activity(OA), linear dichroism(LD), and circular dichroism(CD) — describing the four optical properties of matter.
The potentially complex in Eq. 2 is generally non-diagonalizable in 7 out of 11 cases, unless is a normal matrix[107], corresponding to the 4 possible combinations, that is, OA + Bi, LD + CD, Bi + LD, OA + CD, i.e. two endpoints of the M-M TC’s 4 edges ‘H’, ‘!H’, ‘S’, ‘!S’, respectively.
Leveraging the reciprocal-space M-M TC in Fig. S8, we analyze the pairwise competitions among CD, OA, and LD, and how it affects of the eigensystem pairs’ distribution in 2D domain in Supplementary Note 7. We established a one-to-one extension of the three principal concepts proposed by Berry & Dennis[5] — optical singularities(i.e. EPs or singular axes), L lines, and C points — by (1) generalizing the finite set of eight optical singularities into infinite families with disk-like, annular, and crescent geometries; (2) extending L lines into L lakes and their contours, namely the L shorelines; and (3) broadening the haunting theorem associated with C points from a narrow interpretation (restricted to varing LD) to a general one (also valid for varing CD). We formulated criteria for the equivalences CD OA, OA LD, and CD LD. Ultimately, all of these confirm that eigensystems of any material are computable under our LFCO framework, foreshadowing general field reconstructions in 3D domain.
Having mastered the computation of eigensystems within arbitrary materials, the next task is to batch-reproduce camera-recorded data in real space. For this purpose, both the in-crystal CO eigenmodes with their 32 transition matrix (Eqs. 5 and 6) in reciprocal space and specific out-of-crystal FO OTFs are essential.
We thus design a crystal-2f system in Fig. S9, placing a single lens at 1f(one focal) length from the rear face of the crystal, thereby performing a 2D FT between the two focal planes of the lens. The output field at the crystal’s back surface — calculated via the 32 transition matrix — is projected onto the XY plane at 2f, representing the strict far-field limit of the Fraunhofer diffraction.
To quantitatively validate the proposed crystal-2f system, we scanned the pump’s off-axis angle and the propagation distance in Fig. S9, thereby identifying the precise parameters required to reproduce Peet’s experimental results[198] on internal conical diffraction with Laguerre-Gauss(LG) light beams.
We now revisit the real-space M-M TC in Fig. 7, upon completing the quantitative validation in Fig. S9 around the optic axis, i.e. the diabolic point of a typical biaxial material, where the eigenvalue degeneracy does not extend to the eigenvectors.
Along its three edges, namely ‘H’, ‘S’, and ‘Im’, we broadly reproduce experimental results[113, 239, 217, 247, 159, 158, 139, 81, 160] of Pancharatnam[158], Bloembergen et al.[159], Brenier et al.[139, 81, 160] in 3D real space, focusing on three optical properties of matter, i.e. Bi, OA, and LD.
Perform a counterclockwise scan along the three edges of the M-M TC. First, increasing OA from 0 along the ‘H’ edge of the M-M TC yields the AE diagram of the optical rotation field from vertice ‘Bi’ to ‘OA’ in Fig. 7e (pointed by the light blue arrows), ultimately converging to Bloembergen et al.’s experiment: chiral CR[159] through -HIO3. Next, while maintaining OA, as LD increases along the ‘Im’ edge from 0 (denoted by the light green arrows), Brenier’s experimental results on chirality versus dichroism[160] of acentric chiral Nd3+-doped BZBO are obtained, with a clear criterion for OA LD. That is, when pump’s polarization is aligned parallel to the polarization of Voigt wave while the analyzer is set perpendicular, central extinction is observed in the far field. Following this, by canceling OA and keeping only LD, the angular absorption distribution of laser crystal KGd(WO4)2 is obtained[139, 81]. Then, by gradually reducing LD along the ‘S’ edge (antiparallel to the light red arrows), the Pancharatnam phenomenon[158] emerges.
Finally, as LD is reduced to 0, only Bi remains, including uniaxiality (Fig. S10b and Fig. S11), biaxiality (Fig. S10a and Fig. S11b1), and their hyperbolic counterparts (Fig. S10a,b). As an effect caused by the off-diagonal elements of the 22 eigen-polarization matrix and an embodiment of conservation law, the global grasp and detailed calculation of SOI play a significant role in not only LCO[215, 216, 217, 219, 221, 223, 222, 224, 225, 226] (see Fig. 6 and Figs. S9, S11 and SA3-SA5) but also NCO[247, 281, 6, 256] (see Fig. 4).
2.5 Tightly focused light in highly anisotropic materials
In the extreme case of strong linear interaction between highly anisotropic materials and tightly focused light fields — an area at the forefront of both industry and academia in laser processing, aberration correction, and inverse focal engineering — our LFCO model provides a unified solution, as shown in Fig. 8. Here, we show the underlying mechanism of Raman spike[161] (Fig. 8b) in CR, the reverse engineering technique for the targeted vector complex field at the focal plane free of Zernike polynomial (Fig. 8c1), the “Cherenkov cone” outside the inner beam (Fig. 8c2) marking the computational boundary, the R-L and o-e decomposition analysis for high-N.A. pump (Fig. 8c4), and double conical refraction and the optical field self-twisting effects it induces (Fig. 8d1).
The interference of complex fields gives rise to the intricate and distorted shapes of light spots, especially in the case of tightly focusing. This phenomenon, emblematic of wave optics (paralleling quantum mechanics), as its most distinctive feature differentiating ray optics (analogous to classical mechanics), not only represents the mathematical frontier of highly oscillatory partial differential equations(PDEs), where both the inner integrand and the integral outcome are potentially highly oscillatory[282], but also pushes the computational limits in raising the speed-accuracy product under the constraints of the Nyquist(-Shannon) sampling theorem[283], thus stands as one of the most promising candidates for fully challenging the performance of all neural networks claiming that they are physical[284].
The strong anisotropy of the material further heightens the computational demands for calculating the light field distribution. This arises from the enhanced distortion of the electromagnetic field’s eigenvalue surfaces and their associated wavefronts (equiphase surfaces), compounded by the faster variation of eigenvectors controlling polarization directions which ultimately implement interference along x, y, z axes, hastening the premature arrival of the nightmare where phase differences exceeding between adjacent pixels[66, 68] on a typical planar interference pattern at .
Using the crytal-version vector angular spectrum method(ASM), i.e. this LFCO model, without any optimization, we explored various tightly focused field distributions while operating at the edge of the sampling theorem, with each subfigure (except Fig. 8c2) in Fig. 8 nearing the computational limits just before aliasing errors occur. Initially, we believed that the origin of the outer cone in Fig. 8c2 was physical because: (1) it began forming even before propagation, and (2) it appeared at the center of the image rather than being reflected from the edges. However, we later discovered that it was still caused by aliasing errors originating from circular convolution. This conclusion was drawn because, in order to shift the focal plane of Fig. 8c2 inside the LN materiral, we set the incident field distribution at the material’s entrance to a LP Gaussian beam with a N.A. close to 1 (see the top row of Fig. 8c3), propagating backward by mm. However, when was set to zero, the outer cone disappeared. In contrast, Fig. 8c4 uses the transfer function of the objective lens instead of backward propagation to shift the focal plane below the upper surface of LN, completely eliminating the aliasing error.
We find that by setting negative, thereby transforming the refractive index surface from an ellipsoid to a dielectric-type two-sheet hyperboloid[114, 118, 119, 113, 122], the two optical axes can be brought closer together, as shown in Fig. 8d2. For KTP crystal at 1064 nm, the angle between the two optical axes decreases from 34.6 in the ellipsoidal case to 7.7 in the hyperboloidal case, allowing us to compute the full evolution inside and outside the crystal in a non-paraxial manner over relatively longer propagation distances of the entire crystal-2f system. Optical field knotting in the 1st and 3rd column of Fig. 8d1 are beyond our present mathematical understanding and techniques for explanation.
2.6 Superstructure of this LCO model: NCO QNCO
Nonlinear optics(NO), as a profound gateway for exploring and understanding light-matter interactions, introduces all higher-order nonlinear terms in the CRs beyond the first-order linear electric susceptibility . These higher-order terms represent the nonlinear response of bound electric dipoles to external optical fields. Ultimately, they act as cross-band light sources that coherently generate NFs within crystals through parametric radiation (there are also non-parametric/inelastic cases where phonons or molecules are involved[22, 286]).
Both the NFs and the pumps involved in the interaction, are constrained by their own monochromatic passive LCO wave equation, necessitating independent diffraction as the crystal’s eigenmodes (Fig. 5d,e,g,h).
The NCO parametric frequency conversion process must first satisfy energy conservation, followed by momentum conservation (often described as wave vector or phase matching), which further tests the precision of LCO eigenvalue calculations (Fig. 5d,e). If the anisotropy of the second-order nonlinear coefficient tensor is additionally involved, the accurate computation of eigenvectors (Fig. 5g,h) in the frame[277, 13, 3, 4] within LCO must also be ensured as a prerequisite.
These two core principles, i.e, passive independent diffraction and active coupled conversion of all that participate in NCO processes, anchor all NCO phenomena within the framework of LCO. The precise theoretical modeling of NCO processes in anisotropic materials becomes naturally, a more rigorous test for all established LCO models.
Having conducted large-scale numerical validation of known NCO phenomena in Section 2.2, thereby indirectly substantiating the LCO model advanced in this paper, it is imperative to further establish the model’s forward-looking predictive capabilities, for the potentially vast range of yet-unknown NCO phenomena.
To this purpose, we propose simulations of spontaneous parametric down-conversion(SPDC) in 1D, 2D, and 3D NPCs in Fig. 9 (and Fig. SB6), as a preliminary step toward developing a comprehensive theoretical and mathematical FO framework for quantum NCO(QNCO) nonlinear quantum optics(NQO) QO.
3 Discussion
In arbitrary , namely birefringent-chiral-dichroic dielectrics, complex eigenmodes and their anisotropic diffraction behaviors form the fundamental basis of all advanced studies involving the interaction between light and matter. Of all the CO models created to address this issue, the ultimate plane wave solution in the form of matrix exponential after Jordan decomposition is regarded as the most likely key to deciphering the internal structure of singularities(EPs) with second-order or higher degeneracy (see Supplementary Note 1). Yet, the batch numerical implementation of this approach is unstable and incapable of managing thicker non-Hermitian materials.
Conversely, numerically stable models that allow for the treatment of thick non-Hermitian slabs fail entirely to resolve EPs. These models share the same structure of classical plane wave , whose two distinct forms based on real spherical/rectangular coordinates, with different independent variables , both satisfy the same wave/characteristic Eq. 2. All models choosing as input variable typically requires numerical solutions for quartics, while the rest models depends on possess simple closed-form solutions to biquadratics. However, the former satisfies FT and boundary conditions, while the latter does not.
In light of complex ray tracing, we extend Berry & Dennis’s 2003 uniform plane wave LCO model to non-uniform LFCO and ultimately derive the explicit form of its 32 transition matrix field between any two sections within a planar slab dielectric, by transforming Berry & Dennis’s eigensystem into , thus bridging the two major branches of LCO in reciprocal space, where either ray direction or spatial frequency serves as the input variable.
Using this LFCO model, we have comprehensively explored and revealed a new facet of the LCO from two perspectives: the AE of eigensystems in 2D reciprocal space under the dual competition among the four material properties: linear/circular birefringence/dichroism (via the M-M TC), together with the corresponding evolution of the light field in 3D real space (in collaboration with OTFs exemplified by the crystal-2f setup), depicting two magnificent panoramic maps in both real and reciprocal space.
Along the way, we observe that circular dichroism(CD), which also conforms to the haunting theorem pertaining to C points, can result in heart-formed L shorelines and infinite singularities arranged in disk-, ring-, and crescent-like shapes in 2D domain, implementing a bijective extension of the classical CO theory concerning singular axes, C points, and L lines; while in 3D space, the genesis of Raman spikes during CR, double conical refraction(DCR) and optical knots(OKs) are discovered.
By integrating the custom-developed FO transfer functions of optical instruments, our trilogy (this LFCO model subsequent NFCO models) has also successfully reproduced and further explored numerous complicated experimental results in both LCO and NCO, unveiling the unified mathematical and physical essence underlying these diverse phenomena. This forward LFCO model is also inversely applied for inverse focal design in high N.A. laser writing. Lastly, we offer some initial demonstrations of the potential future applications of this powerful LCO model in (quantum) NCO, including spontaneous parametric down-conversion(SPDC) in 3D nonlinear photonic crystals(NPCs).
Such a novel paradigm (the M-M TC and FO OTFs with crystal-2f configuration) and the newly predicted - and -space phenomena will inject theoretical, experimental and computational vitality into LCO & NCO, as well as CO, FO, LO, NO, and QO, breathing new life into their future developments. It also lays a solid foundation for all upper-level architectures based on light-matter interactions.
In addition to developing NCO models based on this LCO framework, we are also actively exploring new solutions to the LCO model itself. This includes investigating alternative (non-rectangular, non-spherical) coordinate systems or retaining Berry & Dennis’s “South-Pole Stereographic Projection” coordinate system while utilizing non-uniform FT[111] of type 2 (and 1) — sampling non-uniform grid in 2D reciprocal space while uniformly sampling grid in 2D real space to avoid solving transcendental equations.
As an advanced version of linear (wave) optics(LO) in isotropic free space (culminating in classical FO in air or vacuum) — an area older than CO itself (in fact, LO and CO have each developed for over 200 years), LFCO (together with the NFCO derived from it) inherits all of its challenges and extensions. Examples include introducing the (Chriped) Z transform[66], Bluestein algorithm[67], scaled angular spectrum[68], and semi-analytical FT[69] into (L/N)FCO, etc., enabling adaptive field-of-view(FOV) or region-of-interest(ROI) scaling in real and reciprocal spaces as a function of propagation distance , delaying the onset of aliasing errors, and reducing computational complexity in both time and space, all while satisfying the sampling theorem.
It is worth emphasizing that similar problems remain at the forefront of the intersection between applied mathematics and computer science (i.e. numerical analysis). That is, achieving fast and accurate semi-numerical implementation to PDEs of highly oscillatory fields and their integral solutions, even in the one-dimensional ordinary differential case, remains a relentless challenge and an enduring pursuit[282].
As a longstanding issue spanning mathematics (linear algebra, PDEs, fractional FT[288, 289, 63, 290, 291], topology), physics (Hamilton’s five great legacies: manifolds, complex numbers, diabolic points, Hamiltonians, quaternions as a whole, Noether’s theorem, and wave optics — the mother of quantum mechanics), and computer science (sampling theorem, butterfly algorithm, Z transform, physical information and convolutional neural networks), along with a scientific “toy gallery”(see Supplementary Note 9) where a simple laser pointer and a small crystal suffice for real-time experiments to verify surrounding objective reality, we hope the “old yet new” aspects of (L/N)(F)CO will once again attract attention across diverse domains.
4 Methods
4.1 Boundary conditions for laboratory settings
For a commonly used homogeneous dielectric planar slab in laboratory settings, the space dependence of the major material quantity (see Supplementary Note 2) is characterized by two step functions[191, 292, 90], assuming that the slab is surrounded by isotropic non-chiral transparent media. After establishing the 3D laboratory coordinate system (LCS), also referred to as the frame, whose +z-axis is aligned parallel to the inward normal of the front face of the slab with air as its surroundings, the dielectric tensor of the slab is expressed as[191, 292]
Then the wave equation (see Supplementary Note 2), together with Eq. 6, contains all the necessary information for solving the distribution of electromagnetic field inside and outside the slab made of typical optical materials, provided that a two-dimensional(2D) distribution of the vector pump (’s transverse components ) in front of the slab () is given, which means neither divergence equations[191, 292, 231, 240, 91, 90, 293, 294, 74, 75, 76, 77, 78, 295] nor boundary conditions[191, 292] are even required.
If we must assert which boundary conditions are the most fundamental, we would choose the tangential continuity of the electric field and the generalized Snell’s law[86, 84] as the sole boundary conditions, due to the general failure of tangential continuity for the magnetic field [191, 292, 195, 296, 297] and the inconvenient use of the normal continuity for the magnetic induction field as its substitute.
The trade-off is that, for each incident field, either two transmission fields passing through the anti-reflective(AR) coating/nanostructure or two reflected fields bounced back[298, 299] by the high-reflective(HR) coating can be calculated relatively accurately, while finer effects such as the photon spin-Hall effect[300, 301] and the additional lateral shift[243] cannot be revealed by this model.
Above boundary conditions adopted in our model naturally align with the standard configuration in modern nonlinear photonics laboratories where slab-shaped crystals with AR/HR coatings/nanostructures applied to both front and rear surfaces are frequently used, rendering our model applicable in the majority of cases.
Besides, regarding the central element of this work, i.e., the 32 transition matrix of non-uniform LFCO, since it is defined solely within the material, no extra boundary conditions are required from this standpoint, apart from the generalized Snell’s law, which mainly restrains the eigenvalues.
4.2 Eigensystem corrections for phase continuity
The generalized Snell’s law between transparent and dissipative/active media with non-Hermitian , will be violated, if one persists in employing uniform plane waves with uniform complex wave vectors[5, 20, 235, 74, 75, 76, 77, 78, 79, 80, 81]
across the slab material (). Because when sticking to this form, transverse wave vectors are generally complex and -dispersive on the material side (with and ), whereas in air surroundings (where and or ) and FT they remain real and non--dispersive, breaking the in-plane momentum conservation = phase continuity, one-to-one correspondence for each plane wave across interfaces and the requirement of FT.
By contrast, adhering to non-uniform complex wave vectors
of the non-uniform plane waves throughout naturally align with the generalized Snell’s law[86, 84, 14] on the interface planes at and with non-absorbing surroundings. Because through gluing to this form, the entire space () consistently takes -dispersion-free spatial frequencies as transverse wave vectors, which further permit one-to-one correspondence across boundaries for each Fourier component = spatiotemporal spectrum while fulfilling the constraints of FT.
Our approach to amending the complex eigenvalues in the aforementioned complex wave vectors , while retaining their analytical form, is quite straightforward (see Eq. (S12) in Supplementary Note 2):
where both the input real ray direction for uniform eigenvalue and the output non-uniform eigenvalue are ultimately functions of , and mutually coupled, as illustrated in Fig. 10a2.
One can unravel this coupling by first discarding the imaginary part of the complex non-uniform wave vector in Eq. 8b, retaining only the real part ’s unit vector from Eq. 9a as the ray direction in space. The input variable is then inserted into Eq. 9b to form a transcendental equation which asymptotically converges to the ground true eigenvalue of the booker quartic on condition that
where material absorption is sufficiently low with typically when a transmission spectrum is still present, as exemplified by Brenier’s laser crystal[139, 81] in Fig. 7b,e.
Berry & Dennis’s input variable can be obtained from Fig. 10a2 as an intermediate data. Inserting it into Fig. 10b yields Berry & Dennis’s eigensystem , where the eigenvector also converges to the ground truth under the condition of Eq. 10 (see Supplementary Note 2.2), after undergoing the operation shown in Fig. 10c.
4.3 Overall flowchart of this LFCO model
Starting from the atomic inputs, we built the comprehensive workflow diagram in Fig. 11 from the bottom up, referencing three submodules from Fig. 10a,c. All elements in the flowcharts (Figs. 6, 10 and 11) of this article are accompanied by step-by-step derivations and detailed explanations provided in the Supplementary Notes.
Tables 1 and 2 briefly presents the categorical meanings of the graphical elements (nodes and arrows) as ‘CLASSes’ in Figs. 6, 10 and 11. The physical/mathematical significance of the specific instantiated ‘OBJECTs’ within each ‘CLASS’ can be found in Supplementary Notes 2-5.
We thank Sir M.V. Berry for the insights shared through both our online and offline exchanges.
This work was supported by the Natural Science Foundation of Jiangsu Province (BK20240005) and the Fundamental Research Funds for the Central Universities (021314380268).
C.X. conceived, led, and built the entire trilogy starting from empty-nothing (down to atom, up to soul) independently, with his unique long-chain/tree/web (non)linear logic, interspersed with leap-like flashes of divine apocalypto, right before the AI era and his own liberation from VPN restrictions. Y.Z. warmly approved, consistently supported, and responsibly supervised the whole trilogy.
All data generated by this trilogy, along with the parameter file for reconstructing any (sub)figure, can be requested from the first author.